21 April 2015: Tuesday


  • The Giraffe migration broke the delivery of feed files to giraffe last night. ┬áSpent two hours (9pm-11pm) working with Kent and Mike to get it fixed.
  • Mike updated the LDAP ACLs because of the giraffe migration and a failure to test.
  • No other issues in the LDAP update last night (well, other than there was no Banner feed because of the giraffe migration).
  • No issues in this morning’s ACCOUNT update (only one item to process anyway).


  • Clones started.
  • New clients that have issues (I love this game of Russian Roulette – NOT!) after the restart of NetWorker (and upgrade to yesterday.
  • Last couple of clients that would not let ozzie back them up are now fixed and backups have been taken.


  • SHA-2 cert installed to shut Chrome up.


Leave a Reply

You must be logged in to post a comment.